# Assessing uncertainty in Gaussian mixtures-based entropy estimation

### v1

@@ 6.1 Gold prices data
Figure 5: Gaussian mixtures-based entropy estimates for each year from 2014 to 20122, with the associated 95% centered percentile intervals obtained using various bootstrap methods (BS = nonparametric; PB = parametric; WLB = weigthed likelihood bootstrap with weights generated from a Dirichlet with $\alpha=1$ and $\alpha=0.8137$).
@@ 6.2 NBA 2022-23 net rating
The joint distribution of average net rating and net rating entropy for NBA teams in the 2022-23 regular season is shown in Figure 9. The average and entropy values are plotted along the x-axis and y-axis, respectively, with horizontal and vertical bars indicating the 95% WLB credible intervals. The net rating average describes the overall quality and performance level of an NBA team, with higher values indicating a better team. On the other hand, the net rating entropy quantifies the consistency of a team’s performance across the season. Higher entropy values suggest a less constant level of play, with more fluctuation in performance from game to game.
Ideally, teams should aim to have both a large positive average net rating, implying they outscored opponents by a considerable margin over the season, as well as a low net rating entropy, indicating they played at a steady, consistent level throughout the year. The most successful teams are typically those that can maintain an excellent level of play while also minimizing fluctuations and inconsistencies over the season.
Figure 9: Scatterplot of net rating (NetRtg) average and entropy for NBA teams on the 2022-23 regular season, with corresponding 95% credible intervals obtained using the WLB approach with $\alpha=0.8137$.
@@ Appendix A Simulation results
Table 3: Simulation results for data generated from a Gaussian distribution with $\mu=0$ and $\sigma=1$. For each bootstrap method and sample size, the average estimate of entropy, the bootstrap estimates of bias and standard error, and the associated bootstrap percentile intervals are reported.
@@
Figure 10: Simulation results for data generated from a Gaussian distribution with $\mu=0$ and $\sigma=1$. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weigthed likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.
Table 4: Simulation results for data generated from a Mixed Gaussian distribution with $\mu=2$ and $\sigma=1$. For each bootstrap method and sample size, the average estimate of entropy, the bootstrap estimates of bias and standard error, and the associated bootstrap percentile intervals are reported.
@@
Figure 11: Simulation results for data generated from a Mixed Gaussian distribution with $\mu=2$ and $\sigma=1$. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weigthed likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.
Table 5: Simulation results for data generated from a Laplace distribution with $\mu=0$ and $\beta=2$. For each bootstrap method and sample size, the average estimate of entropy, the bootstrap estimates of bias and standard error, and the associated bootstrap percentile intervals are reported.
@@
Figure 12: Simulation results for data generated from a Laplace distribution with $\mu=0$ and $\beta=2$. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weigthed likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.
Table 6: Simulation results for data generated from a bivariate Gaussian distribution with $\mu=[0,0]{}^{\!\top}$ and ${\boldsymbol{\Sigma}=\bigl{[}\begin{smallmatrix}1.0&0.8\\ 0.8&2.0\end{smallmatrix}\bigr{]}}$. For each bootstrap method and sample size, the average estimate of entropy, the bootstrap estimates of bias and standard error, and the associated bootstrap percentile intervals are reported.
@@
Figure 13: Simulation results for data generated from a bivariate Gaussian distribution with $\mu=[0,0]{}^{\!\top}$ and ${\boldsymbol{\Sigma}=\bigl{[}\begin{smallmatrix}1.0&0.8\\ 0.8&2.0\end{smallmatrix}\bigr{]}}$. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weigthed likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.
Table 7: Simulation results for data generated from 10-dimensional independent $\chi^{2}$ distributions with $\text{df}=5$ degrees of freedom each. For each bootstrap method and sample size, the average estimate of entropy, the bootstrap estimates of bias and standard error, and the associated bootstrap percentile intervals are reported.
@@
Figure 14: Simulation results for data generated from 10-dimensional independent $\chi^{2}$ distributions with $\text{df}=5$ degrees of freedom each. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weigthed likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.

### v2

@@ 6.1 Gold prices data
Figure 5: Gaussian mixtures-based entropy estimates for each year from 2014 to 2022, with the associated 95% centered percentile intervals obtained using various bootstrap methods (BS = nonparametric; PB = parametric; WLB = weighted likelihood bootstrap with weights generated from a Dirichlet with $\alpha=1$ and $\alpha=0.8137$).
@@ 6.2 NBA 2022-23 net rating
Overall, for a comprehensive assessment of a team’s performance, it is important to consider both the average net rating and its entropy. The net rating average describes the overall quality and performance level of an NBA team, with higher values indicating stronger teams. In contrast, net rating entropy measures the consistency of a team’s performance across the season, with higher entropy values indicating greater fluctuation from game to game. Ideally, successful teams should aim for both a high positive net rating, reflecting strong overall performance, and low entropy, indicating consistent play.
@@ Appendix A Simulation results
Table 3: Simulation results for data generated from a Gaussian distribution with $\mu=0$ and $\sigma=1$. For each bootstrap method and sample size, the table reports averaged values of the entropy estimates, bootstrap estimates of bias, standard error, and percentile interval limits, along with their empirical coverage.
@@
Figure 9: Simulation results for data generated from a Gaussian distribution with $\mu=0$ and $\sigma=1$. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weighted likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.
Table 4: Simulation results for data generated from a Mixed Gaussian distribution with $\mu=2$ and $\sigma=1$. For each bootstrap method and sample size, the table reports averaged values of the entropy estimates, bootstrap estimates of bias, standard error, and percentile interval limits, along with their empirical coverage.
@@
Figure 10: Simulation results for data generated from a Mixed Gaussian distribution with $\mu=2$ and $\sigma=1$. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weighted likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.
Table 5: Simulation results for data generated from a Laplace distribution with $\mu=0$ and $\beta=2$. For each bootstrap method and sample size, the table reports averaged values of the entropy estimates, bootstrap estimates of bias, standard error, and percentile interval limits, along with their empirical coverage.
@@
Figure 11: Simulation results for data generated from a Laplace distribution with $\mu=0$ and $\beta=2$. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weighted likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.
Table 6: Simulation results for data generated from a bivariate Gaussian distribution with $\mu=[0,0]{}^{\!\top}$ and ${\boldsymbol{\Sigma}=\bigl{[}\begin{smallmatrix}1.0&0.8\\ 0.8&2.0\end{smallmatrix}\bigr{]}}$. For each bootstrap method and sample size, the table reports averaged values of the entropy estimates, bootstrap estimates of bias, standard error, and percentile interval limits, along with their empirical coverage.
@@
Figure 12: Simulation results for data generated from a bivariate Gaussian distribution with $\mu=[0,0]{}^{\!\top}$ and ${\boldsymbol{\Sigma}=\bigl{[}\begin{smallmatrix}1.0&0.8\\ 0.8&2.0\end{smallmatrix}\bigr{]}}$. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weighted likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.
Table 7: Simulation results for data generated from 10-dimensional independent $\chi^{2}$ distributions with $\text{df}=5$ degrees of freedom each. For each bootstrap method and sample size, the table reports averaged values of the entropy estimates, bootstrap estimates of bias, standard error, and percentile interval limits, along with their empirical coverage.
@@
Figure 13: Simulation results for data generated from 10-dimensional independent $\chi^{2}$ distributions with $\text{df}=5$ degrees of freedom each. Panel (a) shows the empirical coverage for 1000 replications as function of the sample size for different bootstrap methods (BS = nonparametric; PB = parametric; WLB = weighted likelihood bootstrap with weights generated from a Dirichlet with different values of $\alpha$) and two type of bootstrap intervals (perc = percentile; cperc = centered percentile). Panel (b) show the ratio of bias (in absolute value) over the standard error for each bootstrap procedure as function of sample size.